OpenStudy (anonymous):

write an equation for a linear function f that has the given values f(-4)=-8, f(-8)=-11?

OpenStudy (hoblos):

the line passes by (-4,-8) &(-8,-11) slope= 3/4 y+8=3/4(x+4) y= 3x/4 -5

OpenStudy (hoblos):

f(x)= 3x/4 -5